### v1

@@ Appendix A Section title of first appendix
1:Initialize: $\mathcal{S}=[s_{1},s_{2},...,s_{m}];\ \mathcal{D}=[d_{1},d_{2},...,d_{m}];\ \mathcal{B}=[b_{1},b_{2},...,b_{m}];\ \mathcal{X}=[x_{1},x_{2},...,x_{n}];\ \mathcal{Y}=[y_{1},y_{2},...,y_{n}]$
@@
4:     $\mathcal{D}\leftarrow\mathcal{D}_{best}$
@@
8:              The nodes: $\mathcal{A}=[a_{1},a_{2},...,a_{m}]$
9:              Constructing a neural network: $\mathcal{A}\leftarrow\mathcal{S}$
@@
11:                  $\mathcal{A}[k]\leftarrow\mathcal{S}[k]$
@@
13:              $\mathcal{E}=[\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{m}]$
@@
15:                  $\mathcal{E}[k]\leftarrow d_{k}*s_{k}(x_{i})+b_{k}$
@@
17:              $\hat{y}\leftarrow\mathcal{E}[0]$
18:              loss $\leftarrow\frac{1}{2}(y_{i}-\hat{y})^{2}$
19:              $\mathcal{E}_{all}=f(X)$
20:              $\mathcal{R}^{2}=r2(\mathcal{E}_{all},y)$
@@
22:              Train($key,\mathcal{E},\mathcal{D},\mathcal{B}$)
@@
Appendix B
@@
4:     $\mathcal{R}_{\text{best}}\leftarrow R^{2}$
5:     $\mathcal{S}_{\text{best}}\leftarrow\mathcal{S}$
6:     $\mathcal{D}_{\text{best}}\leftarrow\mathcal{D}$
7:     $\mathcal{B}_{\text{best}}\leftarrow\mathcal{B}$
@@
Appendix C
@@
Variables: $key$; $\mathcal{E}=[\mathcal{E}_{1},\mathcal{E}_{2},...,\mathcal{E}_{m}]$;
      $\mathcal{D}=[d_{1},d_{2},...,d_{m}]$, $\mathcal{B}=[b_{1},b_{2},...,b_{m}]$
if $key\mod ITE\neq 0$ then
     $\mathcal{D}_{new}[k]\leftarrow\mathcal{D}[k]-\alpha\frac{\partial L}{\partial\mathcal{D}[k]}$
@@
if $key\mod ITE=0$ then
@@
         $\mathcal{E}_{new}[k]\leftarrow\mathcal{E}[k]-\alpha\frac{\partial L}{\partial\mathcal{E}[k]}$
@@
     $\mathcal{E}$)
@@
Appendix D
@@
1:Variables: $key$; $\mathcal{E}_{new}=[\mathcal{E}_{1new},\mathcal{E}_{2new},...,\mathcal{E}_{mnew}]$;
2:$activation\ functions=[+,-,*,/,sin,cos,sqrt,log,exp,sigmoid,relu,x]$
@@
4:     $\mathcal{SY}\leftarrow\{\mathcal{E}[k]_{newl}+\mathcal{E}[k]_{newr},\mathcal{E}[k]_{newl}-\mathcal{E}[k]_{newr},$
5:           $\mathcal{E}[k]_{newl}*\mathcal{E}[k]_{newr},\mathcal{E}[k]_{newl}/\mathcal{E}[k]_{newr},$
6:           $sin(\mathcal{E}[k]_{newl}),cos(\mathcal{E}[k]_{newl}),sqrt(\mathcal{E}[k]_{newl}),$
7:           $log(\mathcal{E}[k]_{newl}),exp(\mathcal{E}[k]_{newl}),sigmoid(\mathcal{E}[k]_{newl}),$
8:           $relu(\mathcal{E}[k]_{newl}),x\}$
9:     $Choice\leftarrow|\mathcal{SY}-\mathcal{E}[k]_{new}|$
10:     $Index\leftarrow\operatorname{arg\,min}(Choice)$
11:     $S[k]\leftarrow activation\ functions[Index]$
@@
Appendix E
@@
Appendix F
(a) Nguyen-1
(b) Nguyen-2
(c) Nguyen-3
(d) Nguyen-4
(e) Nguyen-5
(f) Nguyen-6
(g) Nguyen-7
(h) Nguyen-8
Figure 5: This figure shows the extrapolation of the three algorithms on the unary formula in the Nguyen data set in Table 1.
(a) Nguyen-1
(b) Nguyen-2
(c) Nguyen-3
(d) Nguyen-4
(e) Nguyen-5
(f) Nguyen-6
(g) Nguyen-7
(h) Nguyen-8
Figure 6: This figure shows the extrapolation of the three algorithms on the unary formula in the Nguyen data set in Table 1.
(a) Nguyen-9
(b) Nguyen-10
(c) Nguyen-9-X-Net
(d) Nguyen-10-X-Net
(e) Nguyen-9-ANN
(f) Nguyen-10-ANN
(g) Nguyen-9-SVR
(h) Nguyen-10-SVR
Figure 7: This figure shows the test results of our algorithm, ANN, and SVR on the Nguyen-9 and Nguyen-10 in Table 1. To see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, we drew the contour plots after the difference between the predicted surface and the real surface of the three algorithms respectively.
(a) Nguyen-11
(b) Nguyen-12
(c) Nguyen-11-X-Net
(d) Nguyen-12-X-Net
(e) Nguyen-11-ANN
(f) Nguyen-12-ANN
(g) Nguyen-11-SVR
(h) Nguyen-12-SVR
Figure 8: This figure shows the test results of our algorithm, ANN, and SVR on the Nguyen-11 and Nguyen-12 in Table 1. To clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
(a) Nguyen-9
(b) Nguyen-10
(c) Nguyen-9-X-Net
(d) Nguyen-10-X-Net
(e) Nguyen-9-Kronecker NN
(f) Nguyen-10-Kronecker NN
(g) Nguyen-9-Shinchi
(h) Nguyen-10-Shinchi
Figure 9: This figure shows the test results of our algorithm, Kronecker NN, and Shinichi on the Nguyen-9 and Nguyen-10 in Table 1. In order to clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.
(a) Nguyen-11
(b) Nguyen-12
(c) Nguyen-11-X-Net
(d) Nguyen-12-X-Net
(e) Nguyen-11-Kronecker NN
(f) Nguyen-12-Kronecker NN
(g) Nguyen-11-Shinchi
(h) Nguyen-12-Shinchi
Figure 10: This figure shows the test results of our algorithm, Kronecker NN, and Shinichi on the Nguyen-11 and Nguyen-12 in Table 1. In order to clearly see the difference between the predicted surface and the real surface in the range [-2,2] and outside the range, We calculate the difference between the prediction planes of the three algorithms and the actual plane respectively, then plot the contour lines.

### v2

@@ 2 Pseudocode for X-Net
1:Initialize: $\mathcal{S}=[s_{1},s_{2},...,s_{m}];\ \mathcal{W}=[w_{1},w_{2},...,w_{m}];\ \mathcal{B}=[b_{1},b_{2},...,b_{m}];\ \mathcal{X}=[x_{1},x_{2},...,x_{n}];\ \mathcal{Y}=[y_{1},y_{2},...,y_{n}]$
@@
4:     $\mathcal{W}\leftarrow\mathcal{W}_{best}$
@@
8:              The nodes: $\mathcal{A}=[a_{0},a_{1},...,a_{m}]$triangleright Initializing nodes.
9:              Constructing a tree-like network: $\mathcal{A}\leftarrow\mathcal{S}$
@@
11:                  $\mathcal{A}_{k}\leftarrow\mathcal{S}_{K}$$\triangleright$ Each node is assigned a specific activation function
@@
13:              $E=[E_{0},E_{1},...,E_{m}]$triangleright Initialize each node output
@@
15:                  $E_{k}\leftarrow w_{k}*S_{k}(x_{l},x_{r})+b_{k}$triangleright Compute the output of each node.
@@
17:              $\hat{y}\leftarrow E_{0}$
18:              loss $\leftarrow\mathcal{L}(y_{i},\hat{y})$triangleright Calculating the loss.
19:              $E_{val}=f(X_{val})$
20:              $\mathcal{R}^{2}=r2(E_{val},\mathcal{Y}_{val})$$\triangleright$ Get the $R^{2}$ of the current expression on the validation set.
@@
22:              Train($key,ITE,E,\mathcal{W},\mathcal{B}$)
@@
3 Pseudocode for SaveBest
This pseudocode shows that in a regression task if the current $R^{2}$ is greater than the previous best $R^{2}$, we save the optimal $R^{2}$, parameters, and activation function groups (in order of network preorder traversal).
@@ 3 Pseudocode for SaveBest
4:     $\mathcal{R}_{\text{best}}\leftarrow R^{2}$triangleright Save the best $R^{2}$.
5:     $\mathcal{S}_{\text{best}}\leftarrow\mathcal{S}$triangleright Save the best group of activation functions.
6:     $\mathcal{W}_{\text{best}}\leftarrow\mathcal{W}$$\triangleright$ Save the best $\mathcal{W}$.
7:     $\mathcal{B}_{\text{best}}\leftarrow\mathcal{B}$$\triangleright$ Save the best $\mathcal{B}$.
@@
4 Pseudocode for Train
This pseudocode shows the detailed training process of X-Net. We use alternating backward transmission to optimize parameters $[W,B]$ and node output $E$ alternately. We use the hyperparameters $key$ and $ITE$ to control how often $[W,B]$ and $E$ are optimized.
@@ 4 Pseudocode for Train
Variables: $key$; $ITE$; $E=[E_{1},E_{2},...,E_{m}]$;
      $\mathcal{W}=[w_{1},w_{2},...,w_{m}]$, $\mathcal{B}=[b_{1},b_{2},...,b_{m}]$
if $key\%ITE\neq 0$ then $\triangleright$ If the condition is not satisfied, train and update the $\mathcal{W}$, $\mathcal{B}$
     $\mathcal{W}_{new}[k]\leftarrow\mathcal{W}[k]-\alpha\frac{\partial L}{\partial\mathcal{W}[k]}$
@@
if $key\%ITE=0$ $\triangleright$ If the condition is not satisfied, train and update the outputs of nodes.
@@
         $E_{new}[k]\leftarrow E[k]-\alpha\frac{\partial L}{\partial E[k]}$
@@
     UpdateSymbols($E$$\triangleright$ Update the activate functions.
     $key$ += 1
@@
5 Pseudocode for SaveBest
This pseudocode shows the process of node activation function selection for X-Net. Specifically, we feed the updated input of each node (the output of the child node) into all the candidate activation functions for numerical calculation to obtain $\mathcal{SY}$. Then the activation function corresponding to the closest value in $\mathcal{SY}$ to the updated output of the current node is selected as the activation function of the current node.
@@ 5 Pseudocode for SaveBest
1:Variables: $E_{new}=[E_{1new},E_{2new},...,E_{mnew}]$;
2:$activation\ functions=[+,-,\times,\div,\sin,\cos,\sqrt{,}\log,\exp,sigmoid,relu,x]$
@@
4:     $\mathcal{SY}\leftarrow\{E[k]_{newl}+E[k]_{newr},E[k]_{newl}-E[k]_{newr},$
5:           $E[k]_{newl}\times E[k]_{newr},E[k]_{newl}\div E[k]_{newr},$
6:           $sin(E[k]_{newl}),cos(E[k]_{newl}),sqrt(E[k]_{newl}),$
7:           $log(E[k]_{newl}),exp(E[k]_{newl}),sigmoid(E[k]_{newl}),$
8:           $relu(E[k]_{newl}),x\}$$\triangleright$ The new input of this node is fed into the candidate activation function for calculation.
9:     $Choice\leftarrow|\mathcal{SY}-E[k]_{new}|$
10:     $Index\leftarrow\operatorname{arg\,min}(Choice)$$\triangleright$ Select the new activation function with that section.
11:     $S[k]\leftarrow activation\ functions[Index]$triangleright Activation function replacement.
@@
6 Details of the Nguyen dataset
We evaluated X-Net and MLP on the Nguyen symbolic regression benchmark suite  (?), comprising twelve benchmark expressions widely recognized and utilized within the symbolic regression field  (?). Each benchmark is defined by a ground truth expression, described in Table 4. The curves of these formulas have both high frequency and low frequency, simple and complex, which can better reflect the fitting ability of the algorithm.
@@ 6 Details of the Nguyen dataset
7 Small Sample learning ability
MLP training often requires a large number of training samples. However, in real life, the cost of sample collection is very high in many cases, and it is difficult for us to obtain many samples. Therefore, the ability of the model to learn with a small number of samples is very important. To test the small sample learning ability of X-Net and MLP. We tested it on the Nguyen dataset. Specifically, we first sample 10,000 points for each test data. We then sample 50,100,500, and 1,000 points from these. These points are then used to train X-Net and MLP, respectively. Whenever the training is complete, we use the remaining points as the test set to test the model. The specific results are shown in table 5.
Table 5: Comparison of Small Sample learning ability.
Baselines X-Net MLP
Points 50 100 500 1000 50 100 500 1000
Train test Train test Train test Train test Train test Train test Train test Train test
Nguyen-1 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-2 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.0$ $1.0$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-3 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.98$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-4 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.98$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-5 $0.99$ $0.91$ $0.99$ $0.98$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.29$ $0.99$ $0.75$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-6 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.25$ $0.99$ $0.86$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-7 $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-8 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-9 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $-1.02$ $0.99$ $-1.08$ $0.99$ $0.29$ $0.99$ $0.99$
Nguyen-10 $1.00$ $1.00$ $1.00$. $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.12$ $0.99$ $0.89$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-11 $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $1.00$ $0.99$ $0.70$ $0.99$ $0.93$ $0.99$ $0.99$ $0.99$ $0.99$
Nguyen-12 $0.99$ $0.97$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.99$ $0.79$ $0.99$ $0.84$ $0.99$ $0.95$ $0.99$ $0.98$
Average 0.995 0.987 0.995 0.994 0.995 0.995 0.995 0.995 0.990 0.588 0.990 0.761 0.990 0.928 0.990 0.989
8 Complementary Experiment Results
We compare X-Net with two methods: Kronecker Neural Networks (Kronecker NNs) and Shinichi (?). The brief descriptions of the two methods are:
  • Shinichi (?). This method generates the probability distribution of the network structure and then optimizes the parameters of the distribution, rather than optimizing the network structure directly.
Table6 displays the experimental results. Table 6 shows the accuracy and network structure’s complexity (number of nodes and parameters) of the three algorithms on regression and classification tasks.
Table 6: A comparative analysis is presented, focusing on the accuracy and complexity of the final network structures across three algorithms in both regression and classification tasks.
REGRESSION
Benchmark X-Net Kronecker NN Shinichi (?)
$R^{2}$ Nod Para $R^{2}$ Nod Para $R^{2}$ Nod Para
Nguyen-1 1.0 5 22 0.99 12 61 0.99 12 52
Nguyen-2 1.0 9 38 0.99 16 97 0.99 15 79
Nguyen-3 1.0 14 58 0.99 14 78 0.99 11 47
Nguyen-4 0.99 20 82 0.99 24 193 0.99 18 115
Nguyen-5 0.99 5 16 0.99 28 253 0.99 30 289
Nguyen-6 0.99 6 18 0.99 8 33 0.99 7 27
Nguyen-7 0.99 5 16 0.99 10 46 0.99 8 37
Nguyen-8 1.0 1 4 0.99 16 97 0.99 12 61
Nguyen-9 1.0 7 24 0.99 28 253 0.99 21 139
Nguyen-10 1.0 4 14 0.99 80 1761 0.99 56 919
Nguyen-11 1.0 3 10 0.99 20 141 0.99 12 65
Nguyen-12 0.99 9 40 0.99 12 61 0.99 14 73
Average 0.996 7.33 28.50 0.990 21.5 256.17 0.990 18 158.58
CLASSION
Benchmark X-Net Kronecker NN Shinichi (?)
Acc Nod Para Acc Nod Para Acc Nod Para
Iris 98.7% 28 112 98.3% 66 798 98.6% 68 823
Mnist(6-dim) 89.4% 65 244 88.8% 276 13328 88.9% 256 12194
Mnist 99.5% 816 3084 99.2% 682 228073 99.0% 658 201291
Fashion-MNIST(6-dim) 76.2% 122 486 75.3% 364 22432 75.2% 344 21276
Fashion-MNIST 94.1% 1066 3884 94.5% 1167 505159 94.3% 1088 448425
CIFAR-10(6-dim) 26.4% 206 764 22.7% 398 24338 25.2% 344 19784
CIFAR-10 46.8% 2733 10072 44.8% 1932 743355 45.1% 1894 726110
Average 75.9% 719.43 2663.71 74.8% 697.86 219640.43 75.2% 497 204271.85
9 X-Net Powers Scientific Discovery
(a) Performance of the Ada-$\alpha$ Activation Function
(b) Boston Housing Price Data Correlation Matrix.
(c) Predicting housing prices using the ’RM’
(d) Predicting housing prices using the ’LSTAT’
(e) Predicting housing prices using ’RM’ and ’LSTAT’
Figure 4: Figure a depicts the change in algorithmic efficiency before and after the use of ada-$\alpha$; Figure b presents the correlation matrix among different variables in the Boston housing price prediction; Figure c demonstrates the schematic representation of housing price predictions using the variable ‘RM’. It can be deduced that the housing prices are directly proportional to the variable ‘RM’, which is consistent with the findings presented in Figure c; Figure d displays the results of predicting housing prices using the variable ‘LSTAT’. Figure e showcases the schematic representation of predicting housing prices using both ‘RM’ and ‘LSTAT’.
9.1 Economic Modeling
The Boston house price forecast data set  (?) is a classic data set in machine learning. Although it is simple, it can test the comprehensive performance of the algorithm well. Each sample in the Boston housing price data set is composed of 12 characteristic variables such as ’CRIM’, ’ZN’, and housing price MEDV. The thermal map of the correlation coefficient between each variable is shown in the figure4b. From the heat map 4b, we can find that the variable ’RM’ (the number of rooms per house) has the highest correlation coefficient with the MEDV(housing price), which is 0.70, showing a positive correlation with the housing price. In contrast, the variable ’LSTAT’ (how many people in the area are classified as low-income) has a minimum correlation coefficient of -0.74 with housing prices, showing a negative correlation with housing prices.
$MEDV=9.10RM-34.67(RM\geq 2)$ (18)
The predicted results are shown in Fig. 4c. From Formula 18, we can find that the final expression obtained by our algorithm can well reflect the positive correlation between RM and housing price.
We used LSTAT as input to predict housing prices using X-Net and finally got formula (19), from which we could see that the formula also well reflected the negative correlation between LSTAT and MEDV. Show in Fig 4d.
$MEDV=\frac{81.39}{(LSTAT+0.44)^{0.44}}-6.54(0\leq LSTAT\leq 100)$ (19)
Finally, we use RM and LSTAT as inputs to predict housing prices, and we get formula (20), the predicted results are shown in Fig. 4e
$MEDV=\frac{15.6(0.48RM-13.21)}{2.36LSTAT-24.53}+9.97$ (20)
(a) Solar energy power generation fitting result
(b) Figure g after sorting by $y_{pre}$
(c) Wind energy power generation fitting result
(d) Figure i after sorting by $y_{pre}$
(e) The correlation matrix in wind energy prediction
Figure 5: Figure a and Figure b show the prediction results of Formula 21 for solar power generation; Figure c and Figure d display the fitting results of Formula 22 for wind power generation data; Figure e represents the correlation coefficient matrix of the variables used in the wind power generation data.
From Formula 20, we can clearly find that the housing price is directly proportional to the variable RM and inversely proportional to the variable LSTAT. That is the more rooms in the house, the more expensive the house, the more low-income people in the community, and the lower the price. The result of this prediction is completely in line with the facts. The above results prove that our algorithm can well reflect the relationship between each variable and the result even for multiple variables.
9.2 Modeling in energy science (Solar And Wind Power Generation Forecasting) (?)
$Y=x_{1}(0.00044x_{3}(-0.00003x_{1}x_{2}+0.31)+0.093)-0.2466$ (21)
The fitted curves of the sun power are shown in Figure 5a and 5b. Note: Figure 5b underwent the same data sorting process as Figure 2b.
The fitted curves of the wind power are shown in Figure 5c and 5d. Note: Figure 5d underwent the same data sorting process as Figure 2b. From Equation 22, we can see that X-Net has modeled the data well using only the first variable (wind speed) among the three variables. This is consistent with our understanding that wind power generation is mainly related to wind speed, which also aligns with reality. Furthermore, from the correlation coefficient heatmap 5e, we can clearly see that the correlation between $x_{1}$ (wind speed) and $y$ (power generation) is as high as 0.86.
10 Ada-$\alpha$ function performance analysis
In the process of neural network training, whether the step size is appropriate directly affects the speed of network convergence. If the step size is excessively large, the optimal solution may be skipped, If the step size is too small, the convergence rate may be extremely sluggish, and the optimal solution may not be obtained within a predetermined number of iterations. Therefore, to improve network efficiency, we design a step-size dynamic adaptive function Ada-$\alpha$. In order to test its performance, we selected four functions ’Nguyen-2’, ’Nguyen-6’, ’Nguyen-9’, and ’Nguyen-10’ from the Nguyen data set to test the Ada-$\alpha$. Using the Ada-$\alpha$ function, we execute each of the formulas ten times, record the amount of time required to fully recover the formula, and then calculate an average. Then do the same for the case without the Ada-$\alpha$ function. The specific bar graph is depicted in Fig. 4a. We can observe that when the Ada-$alpha$ function is implemented, the network’s convergence speed will increase dramatically.